### 06 Mar 2026 v1

# An extension of Birkhoff’s representation theorem to locally-finite distributive lattices

Abstract.
Birkhoff’s representation theorem for finite distributive lattices states that any finite distributive lattice is isomorphic to the lattice of order ideals (lower sets) of the partial order of the join-irreducible elements of the lattice. We present a simplified version of Stone’s extension of this theorem to general distributive lattices. We then apply this formulation to locally finite distributive lattices to produce a novel representation theorem: The lattice is isomorphic to the order ideals of the poset of prime filters of the lattice whose symmetric difference from a particular ideal is finite.
1. Introduction
Birkhoff’s representation theorem for finite distributive lattices has long been known.
Theorem 1 (Birkhoff [Birk1937a][Birk1967a]*§III.3 p. 58).
Any finite distributive lattice $\mathcal{L}$ is isomorphic to the lattice of order ideals of the partial order of the join-irreducible elements of $\mathcal{L}$.
The essence of the proof is every element $x\in\mathcal{L}$ is mapped into the order ideal consisting of the join-irreducible elements of $\mathcal{L}$ that are $\leq x$. The the meet and join of $\mathcal{L}$ correspond to intersection and union of order ideals.
A lattice is finitary if every principal ideal (the set of elements $\leq$ a given element) is finite. Stanley shows that the representation theorem can be extended to finitary distributive lattices.
Theorem 2.
[Stan2012a]*Prop. 3.4.3 Any finitary distributive lattice $\mathcal{L}$ is isomorphic to the lattice of finite order ideals of the partial order of the join-irreducible elements of $\mathcal{L}$.
In combinatorics many interesting lattices aren’t finitary but they are locally-finite, so it is desirable to extend the representation theorem to locally-finite lattices in some way. Two obstacles immediately present themselves. One is that some locally-finite lattices have no join-irreducible elements at all, making it unclear what a representation could be constructed from. The premier example of this is $\mathbb{Z}\times\mathbb{Z}$, as shown in fig. 2.
Figure 1. The lattice $\mathbb{Z}\times\mathbb{Z}$
Figure 2. The lattice $\mathbb{Z}\times\mathbb{Z}$ with some filters outlined
The second obstacle is that in many useful cases, the lattice is isomorphic to the lattice of some, but not all, of the order ideals of join-irreducible elements, all of which are infinite. Often these ideals consist of all the ideals with the exception of the empty ideal and the ideal of all join-irreducible elements. A typical example is shown in fig. 3.
Figure 3. The lattice $t$-SkewStrip (right) and its poset of join-irreducible elements $t$-Plait (left). (Taken from [Fom1994a]*Fig. 7.)
Thus our major goal is a general theory of representation of locally-finite distributive lattices with a secondary goal of understanding lattices that are isomorphic to the lattice of some, but not all, order ideals of its join-irreducible elements.
2. Representation of non-finitary lattices
Our analysis of general distributive lattices follows the discussion in [Nat2017a] and our references to it refer to chapter 8 if not otherwise specified. [^1] Prof. Nation informs me that the discussion in [Nat2017a] is derived from Bob Dilworth’s lattice theory course at Caltech in 1970–71.
This discussion is an alternative to Stone’s presentation of a representation [Stone1938a] that avoids using the machinery of topology, which makes its application to discrete lattices clearer.
For the remainder of this article, let $\mathcal{L}$ be a distributive lattice. As an example, consider $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 2
Definition 3.
Let $P$ be a poset. We define that $F\subset P$ is an order filter (of $P$) iff $x\in P$ and $x\leq z$ imply $z\in F$. We define that $I\subset P$ is an order ideal (of $P$) iff $x\in P$ and $z\leq x$ imply $z\in I$.
Definition 4.
[Nat2017a]*sec. 7 We define that $f\subset\mathcal{L}$ is a lattice filter (of $\mathcal{L}$) iff (1) $f$ is non-empty, (2) $x,y\in f$ implies $x\vee y\in f$, and (3) $x\in f$ and $x\leq z$ imply $z\in f$. We define that $i\subset\mathcal{L}$ is a lattice ideal (of $\mathcal{L}$) iff (1) $i$ is non-empty, (2) $x,y\in i$ implies $x\wedge y\in i$, and (3) $x\in i$ and $z\leq x$ imply $z\in i$.
Definition 5.
[Nat2017a] We define that $f$ is a proper lattice filter (of $P$) iff $f$ is a lattice filter and $f$ is not $\mathcal{L}$. We define a proper lattice ideal dually.
Definition 6.
We define $\mathcal{F}$ to be the poset of lattice filters of $\mathcal{L}$, ordered by inverse inclusion.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, some proper filters are shown in fig. 2 and the corresponding lattice $\mathcal{F}$ is shown in fig. 5.
Figure 4. The lattice of filters $\mathcal{F}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its prime elements outlined, all of which are sur-primes
Figure 5. The poset of prime filters $\mathcal{P}$ for $\mathcal{L}=\mathbb{Z}$. $\mathcal{P}=A\sqcup B$
Definition 7.
Given $f,g\in\mathcal{F}$, we define $f\mathop{\wedge\mkern-8.0mu\wedge}g=\{x\wedge y\,|\,x\in f\textup{ and }y\in g\}$. [^2] We pronounce $f\mathop{\wedge\mkern-8.0mu\wedge}g$ as “f union-meet g” due to lem. 9.
Lemma 8.
Given $\in\mathcal{F}, fqf\mathop{\wedge\mkern-8.0mu\wedge}q$ contains $f$ and $q$.
Proof.
Given any $x\in f$ we can choose any $y\in g$ and then $x\vee y\in g$, so $x=x\wedge(x\vee y)$, showing $x\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Similarly, $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains $g$. ∎
Lemma 9.
$f\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter and is the meet-closure of $f\cup g$.
Proof.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is non-empty: This is trivial because $f$ and $g$ are non-empty.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is meet-closed: Given two elements of $f\mathop{\wedge\mkern-8.0mu\wedge}g$, $x\wedge y$ and $x^{\prime}\wedge y^{\prime}$ with $x,x^{\prime}\in f$ and $y,y^{\prime}\in g$, since $f$ and $g$ are meet-closed, $x\wedge x^{\prime}\in f$ and $y\wedge y^{\prime}\in g$. Then $(x\wedge y)\wedge(x^{\prime}\wedge y^{\prime})=(x\wedge x^{\prime})\wedge(y\wedge y^{\prime})\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter: Given $x\in f$, $y\in g$ and $z\in\mathcal{L}$ with $x\wedge y\leq z$ then $z\vee x\in f$ and $z\vee y\in g$. By distributivity, $(z\vee x)\wedge(z\vee y)=z\vee(x\wedge y)=z$, showing $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$: Trivially, $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains meet-closure of $f\cup g$: Since meet is commutative, any element $z$ in the meet-closure of $f\cup g$ must be the meet of finite non-empty subsets of elements of $f$ and $g$, $z=x_{1}\wedge x_{2}\wedge x_{3}\wedge\cdots\wedge x_{n}\wedge y_{1}\wedge y_{2}\wedge y_{3}\wedge\cdots\wedge y_{m}$ with $n$ and $m$ not both 0, $x_{i}\in f$, and $y_{i}\in g$. If $n=0$, then $z\in g$, and by lem. 8, $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Similarly, if $m=0$, then $z\in f$, and so $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Otherwise, since $f$ and $g$ are meet-closed themselves, $x^{\prime}=x_{1}\wedge x_{2}\wedge x_{3}\wedge\cdots\wedge x_{n}\in f$ and $y^{\prime}=y_{1}\wedge y_{2}\wedge y_{3}\wedge\cdots\wedge y_{m}\in g$, so $z=x^{\prime}\wedge y^{\prime}\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. ∎
Lemma 10.
[…]
Proof.
Regarding join: That the join of two lattice filters $f$ and $g$ is $f\cap g$ is straightforward: $f\cap g$ is necessarily a lattice filter. $f\cap g$ is $\geq$ (that is, is contained in) $f$ and $g$. Clearly, any lattice filter $h\geq f,g$ is contained in both $f$ and $g$, so $h$ must be contained in $f\cap g$, and so must be $\geq f\cap g$.
Regarding meet: Given two lattice filters, $f$ and $g$, let $h=f\mathop{\wedge\mkern-8.0mu\wedge}g$. By lem. 8, $f\subset h$, which is $h\leq f$. Similarly $h\leq g$. To show that $h$ is the greatest lower bound of $f$ and $g$, assume $s$ is an lattice filter and $s\leq f,g$, that is, $f\subset s$ and $g\subset s$. Consider an arbitrary element $z$ of $h$, so $z=x\wedge y$ for some $x\in f$ and $y\in g$. Then $x$ and $y$ are $\in s$, and since $s$ is closed under meets, $z=x\wedge y\in s$. Thus $h\subset s$, which is $s\leq h$. ∎
Lemma 11.
$\mathcal{F}$ is a distributive lattice.
Proof.
Since we know $\mathcal{F}$ is a lattice, it suffices to prove that for any $f,g,h\in\mathcal{F}$, $f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)=(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$, which is equivalent to showing that for any $x\in\mathcal{L}$, $x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$ iff $x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$. The following are equivalent:
$x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$
$x\in f\textup{ and }x\in g\mathop{\wedge\mkern-8.0mu\wedge}h$
$x\in f\textup{ and }(\exists y\in g,z\in h)\,x=y\wedge z$ (1)
To derive (2) from (1): Since $x=y\wedge z=x\vee(y\wedge z)=(x\vee y)\wedge(x\vee z)$ and $f$, $g$, and $h$ are lattice filters,
$x\in f\textup{ and }(\exists y\in g,z\in h)\,x\vee y\in f\cap g\textup{ and }x\vee z\in f\cap h\textup{ and }x=(x\vee y)\wedge(x\vee z)$
Then choose $y^{\prime}=x\vee y$ and $z^{\prime}=x\vee z$. To derive (1) from (2): Choose $y=y^{\prime}$ and $z=z^{\prime}$, and since $f$ is closed under meets. Continuing both cases, (1) is equivalent to:
$(\exists y^{\prime}\in f\cap g,z^{\prime}\in f\cap h)\,x=y^{\prime}\wedge z^{\prime}$ (2)
$x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$
Definition 12.
[Nat2017a] We define that a lattice filter $F$ is prime iff $F$ is proper and if $x\vee y\in F$ implies $x\in F$ or $y\in F$. We define a prime lattice ideal dually.
Lemma 13.
[Nat2017a] A lattice filter $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice ideal iff $\mathcal{L}\setminus f$ is a prime lattice ideal. And dually for lattice ideals.
Theorem 14.
[Nat2017a] For $\in\mathcal{L}, x{{x}\mathclose{\uparrow}}$ is a prime lattice filter iff $x$ is join-irreducible in $\mathcal{L}$.
Note that if $\mathcal{L}$ is not distributive, the condition in this theorem must be modified to “$x$ is join-prime”.
Definition 15.
We define $\mathcal{P}$ to be the poset of prime elements of $\mathcal{F}$, with the order (inverse inclusion) inherited from $\mathcal{F}$.
Definition 16.
For $x\in\mathcal{L}$, we define ${{x}\mathclose{\uparrow}}=\{y\in\mathcal{L}\,|\,x\leq y\}$. For $x\in\mathcal{L}$, we define ${{x}\mathclose{\downarrow}}=\{y\in\mathcal{L}\,|\,y\leq x\}$.
Lemma 17.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$, the following are equivalent:
  1. (1)
    $p\leq{{x}\mathclose{\uparrow}}$,
  2. (2)
    ${{x}\mathclose{\uparrow}}\subset p$, and
  3. (3)
    $x\in p$.
Theorem 18.
[…]
In light of th. 14 we can consider $\mathcal{F}$ to be an extension of $\mathcal{L}$ which “adds the missing join-irreducible elements of $\mathcal{L}$.
Theorem 19.
[Nat2017a]*Th. 8.8 Sublem. A If $x\in\mathcal{L}$ and ${{x}\mathclose{\uparrow}}\not\geq f$ in $\mathcal{F}$, then there exists a $p\in\mathcal{P}$ for which $p\leq f$ but $p\not\leq{{x}\mathclose{\uparrow}}$.
Proof.
Define $\{h\in\mathcal{F}\,|\,h\leq f\textup{ and }h\not\leq{{x}\mathclose{\uparrow}}\}. HH$ is non-empty because $f\in H$. $H$ is a poset ordered by inverse inclusion, inherited from $\mathcal{F}$. Using the Hausdorff maximal principle, select $P$ which is a maximal totally ordered subset of $H$. Since $H$ is non-empty, $P$ is non-empty. Define $p=\bigcup P$. Because $P$ is a non-empty nested set of lattice filters, $\bigcup P$ is a lattice filter also, so $p\in\mathcal{F}$.
Because $x$ is in no member of $P$, $x\not\in\bigcup P=p$ and $p\not\leq{{x}\mathclose{\uparrow}}$.
Since $P$ contains at least one element $q$, $q\leq f$, $f\subset q$, $f\subset\bigcup P=p$, so $p\leq f$.
We prove $p$ is join-irreducible in $\mathcal{F}$: Assume there are $g,h\in\mathcal{F}$ for which $p=g\vee h$ and $g,h<p$ Since $p\not\leq{{x}\mathclose{\uparrow}}$, either $g\not\leq{{x}\mathclose{\uparrow}}$ or $h\not\leq{{x}\mathclose{\uparrow}}$. Without loss of generality, assume $g\not\leq{{x}\mathclose{\uparrow}}$, so $g\in H$.
But since $g<p=\bigcup P$, $\bigcup P\subset g$ and $\bigcup P\neq g$. Those imply that $g$ strictly contains every element of $P$ and $g$ is strictly less than (under inverse inclusion) every element of $P$, and so $g\not\in P$. Thus $P\sqcup\{g\}$ is a totally ordered subset of $\mathcal{F}$ that is strictly larger than $P$, which contradicts that $P$ is maximal. Thus $p$ is join-irreducible. ∎
The Hausdorff maximal principal is equivalent to the axiom of choice.
Corollary 20.
[Nat2017a]*Th. 8.4 Given $x\not\leq y$ in $\mathcal{L}$, there exists a $p\in\mathcal{P}$ for which $x\in p$ and $y\not\in p$.
Definition 21.
An element of $\mathcal{P}$ (which must be prime in $\mathcal{F}$) but is not ${x}\mathclose{\uparrow}$} for any $x\in\mathcal{L}$ (which would have to be join-irreducible) we define to be a sur-prime or sur-irreducible of $\mathcal{F}$ (or of $\mathcal{P}$ or $\mathcal{L}$, by abuse of language).
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, fig. 5 shows that all primes are sur-primes, and not ${x}\mathclose{\uparrow}$} for $x\in\mathcal{L}$. Conversely, added elements need not be sur-primes: $\mathbb{Z}\times\mathbb{Z}$ is in $\mathcal{F}$, not a ${x}\mathclose{\uparrow}$}, but also is not prime.
Theorem 22.
$f\in\mathcal{F}$ is join-irreducible in $\mathcal{F}$ iff $f$ is a prime lattice filter.
Proof.
We will prove the contrapositive: $f\in\mathcal{F}$ is join-reducible in $\mathcal{F}$ iff $f$ is not prime.
Regarding $\Rightarrow$: Assume $f$ is join-reducible. Since $f$ is join-reducible, there exist filters $g$ and $h$ for which $g,h\neq f$ and $f=g\vee h=g\cap h$. Choose $x\in g\setminus f$ and $y\in h\setminus f$; thus $x,y\not\in f$. Since $g$ and $h$ are filters, $x\vee y\in g,h$, and so $x\vee y\in f$. Thus $f$ is not prime.
Regarding $\Leftarrow$: Assuming $f$ is not prime, there exists $x$ and $y$ with $x,y\not\in f$ but $x\vee y\in f$. Define $g={{x}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$ and $h={{y}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}h$, both of which are $\neq f$. We will show that $f=g\vee h=g\cap f$. Clearly $f\subset g\cap h$. Conversely, suppose $z\in g\cap h$. Then there exists $u,u^{\prime},v,v^{\prime}\in f$ for which $z=u\wedge u^{\prime}=v\wedge v^{\prime}$, $u\geq x$, $v\geq y$, and $u^{\prime},v^{\prime}\in f$. Then $z=(u\wedge u^{\prime})\vee(v\wedge v^{\prime})=(u\vee v)\wedge(u\vee v^{\prime})\wedge(u^{\prime}\vee v)\wedge(u^{\prime}\vee v^{\prime})$. But all four terms of that meet are $\in f$, so $z\in f$, showing $g\cap h\subset f$. ∎
Definition 23.
We define $\mathcal{I}$ to be the lattice of order ideals of $\mathcal{P}$ under inclusion, which has $\cap$ as meet and $\cup$ as join.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, the corresponding lattice $\mathcal{I}$ is shown in fig. 6.
Figure 6. The lattice of ideals $\mathcal{I}$ of $\mathcal{P}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its connected components outlined.
Definition 24.
We define $\phi(x)=\{p\in\mathcal{P}\,|\,x\in p\}$.
Lem. 17 can be extended:
Lemma 25.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$, the following are equivalent:
  1. (1)
    $p\leq{{x}\mathclose{\uparrow}}$,
  2. (2)
    ${{x}\mathclose{\uparrow}}\subset p$,
  3. (3)
    $x\in p$, and
  4. (4)
    $p\in\phi(x)$,
Theorem 26.
[Nat2017a]*Th. 8.5 $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$.
Corollary 27.
Every distributive lattice can be represented as a “ring of sets” under intersection and union.
Application to finitary lattices
Theorem 28.
[Birk1937a][Birk1967a]*§III.3 p. 58[Stan2012a]*Prop. 3.4.3 If $\mathcal{L}$ is finitary then every element $p\in\mathcal{F}$ is ${x}\mathclose{\uparrow}$} for some $x\in\mathcal{L}$.
Proof.
Let $p\in\mathcal{F}$. Since $p$ is non-empty and $\mathcal{L}$ is finitary, there must be a minimal element $x\in p$. Since $p$ is closed under meets, every element of $y\in p$ must be $\geq x$, as otherwise $y\wedge x$ would be be <x<x and $\in p$. Thus $x$ is minimum in $p$. Since $p$ is a lattice filter, $p={{x}\mathclose{\uparrow}}$. ∎
Corollary 29.
If $\mathcal{L}$ is finitary then: (1) $\mathcal{P}$ has no sur-primes, so $\mathcal{P}$ is (isomorphic to) the poset of join-irreducible elements of $\mathcal{L}$. (2) The image of $\phi$ is the set of finite elements of $\mathcal{I}$, so $\phi$ is a lattice isomorphism of $\mathcal{L}$ to that set.
Application to locally-finite lattices
In this section, we additionally assume that $\mathcal{L}$ is locally-finite (but not necessarily finitary).
Lemma 30.
If $x\in\mathcal{L}$ is join-irreducible, then $x$ covers exactly one element of $\mathcal{L}$.
Proof.
Since $x$ is join-irreducible, it cannot be the $\hat{0}$ of $\mathcal{L}$ if there is one, so there is an element $y<x$. Since $\mathcal{L}$ is locally-finite, there must be an element $z$ such that $y\leq z\lessdot x$. If $x$ covered more than one element of $\mathcal{L}$, then $x$ would be the join of any two of those elements and would not be join-irreducible. ∎
*** may no longer be needed
Definition 31.
For any $x\in\mathcal{L}$ that is join-irreducible, we define $x^{-}$ (relative to $\mathcal{L}$) to be the unique element that $x$ covers.
Lemma 32.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$ where $x\not\in p$, There exists a $y\in\mathcal{L}$ for which $y>x$, $y\in p$, and $y$ is the minimum element of ${{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$ and thus ${{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$.
Proof.
Define $f={{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$. Since $f$ is a lattice filter, it is not empty, and we can choose $y\in f$, so $y\geq x$. Since $x\not\in f$, $y>x$. And since $\mathcal{L}$ is locally-finite and all choices for $y$ are bounded below by $x$, we can choose $y$ to be minimal in $f$. Any minimal $y$ must be the minimum of $f$, because if $y_{1}\neq y_{2}$ are both minimal in $f$, since $f$ is a lattice filter, $y_{1}\wedge y_{2}$ is also in $f$, contradicting that $y_{1}$ and $y_{2}$ are minimal in $f$. Thus, $f={{y}\mathclose{\uparrow}}$. ∎
Definition 33.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$ where $x\not\in p$, we define $x\mathord{\Uparrow}p$ to be the minimum element of ${{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$.
Dually, we define the parallel operations on lattice ideals:
Lemma 34.
Given $x\in\mathcal{L}$ and $p^{\prime}\in\mathcal{P}^{\prime}$ where $x\not\in p^{\prime}$, There exists a $y\in\mathcal{L}$ for which $y<x$, $y\in p^{\prime}$, and $y$ is the maximum element of ${{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\uparrow}}\cap p^{\prime}$ and thus ${{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\downarrow}}\cap p^{\prime}$.
Proof.
Define $f^{\prime}={{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\downarrow}}\cap p^{\prime}$. Since $f^{\prime}$ is a lattice ideal, it is not empty, and we can choose $y\in f^{\prime}$, so $y\leq x$. Since $x\not\in f^{\prime}$, $y<x$. And since $\mathcal{L}$ is locally-finite and all choices for $y$ are bounded above by $x$, we can choose $y$ to be maximal in $f^{\prime}$. Any maximal $y$ must be the minimum of $f^{\prime}$, because if $y_{1}\neq y_{2}$ are both maximal in $f$, since $f^{\prime}$ is a lattice ideal, $y_{1}\vee y_{2}$ is also in $f^{\prime}$, contradicting that $y_{1}$ and $y_{2}$ are maximal in $f^{\prime}$. Thus, $f^{\prime}={{y}\mathclose{\downarrow}}$. ∎
Definition 35.
Given $x\in\mathcal{L}$ and $p^{\prime}\in\mathcal{P}^{\prime}$ where $x\not\in p^{\prime}$, we define $x\mathord{\Downarrow}p^{\prime}$ to be the maximum element of ${{x}\mathclose{\downarrow}}\wedge p={{x}\mathclose{\downarrow}}\cap p^{\prime}$.
Lemma 36.
If $x\lessdot y$ in $\mathcal{L}$, then $\phi(x)\subset\phi(y)$ and $\phi(y)\setminus\phi(x)$ contains exactly one element.
Proof.
By the definition of $\phi$, it is clear that $\phi(x)\subset\phi(y)$.
By cor. 20, there exists a $p\in\mathcal{P}$ for which $y\in p$ and $x\not\in p$. Thus $p\in\phi(y)\setminus\phi(x)$.
Assume that $p\neq q$ are both $\in\phi(y)\setminus\phi(x)$. Consider $\mathord{\Uparrow}p: x<x$$x\mathord{\Uparrow}p\leq y$. Since $x\lessdot y$, $x\mathord{\Uparrow}p=y$ and $p={{y}\mathclose{\uparrow}}$. Similarly $q={{y}\mathclose{\uparrow}}$, and so $p=q$.
Thus $\phi(y)\setminus\phi(x)$ has exactly one element. ∎
Definition 37.
Given a covering $x\lessdot y$, we define the separator of $x$ and $y$, $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y$ to be unique element of $\mathcal{P}$ that is in $\phi(y)\setminus\phi(x)$. (lem. 36
Lemma 38.
$\mathcal{L}$ is graded. If $x\leq y$ in $\mathcal{L}$, then $\phi(x)\subset\phi(y)$ and $|\phi(y)\setminus\phi(x)|=\rho(x,y)$, where $\rho(x,y)$ is the rank difference of $y$ above $x$.
Lemma 39.
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\not\in\phi(x)$ and $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$. Then $x\lessdot x\mathord{\Uparrow}p$ and $\phi(x\mathord{\Uparrow}p)=\phi(x)\sqcup\{p\}$.
Proof.
Define $y=x\mathord{\Uparrow}p$.
We will show that $\phi(y)=\phi(x)\sqcup\{p\}$. Regarding $\phi(x)\sqcup\{p\}\subset\phi(y)$: Since $x<y$, $\phi(x)\subset\phi(y)$. Since $y\in p$, $p\in\phi(y)$. Thus $\phi(x)\sqcup\{p\}\subset\phi(y)$.
Regarding $\phi(y)\subset\phi(x)\sqcup\{p\}$: For any $q\in\phi(y)$, $q\leq{{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p$. Since $\mathcal{P}$ is distributive and $q$ is join-irreducible in $\mathcal{P}$, $q\leq{{x}\mathclose{\uparrow}}$ or $q\leq p$. That is equivalent to $q\in\phi(x)$ or $q\leq p$. Since $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\in\phi(x)\sqcup\{p\}$.
By lem. 32, $x<y$, and since $\phi(y)$ has exactly one more element than $\phi(x)$, lem. 38 shows that $x\lessdot y$. ∎
Lemma 40.
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\in\phi(x)$ and $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$. Then $x\gtrdot x\mathord{\Downarrow}(\mathcal{L}\setminus p)$ and $\phi(x\mathord{\Downarrow}p)=\phi(x)\setminus\{p\}$.
Proof.
The proof is dual to the proof of lem. 39 but that is concealed by our notation, since $\phi$ is not self-dual.
Define $p^{\prime}=\mathcal{L}\setminus p$ and $y=x\mathord{\Downarrow}p^{\prime}$.
We will show that $\phi(y)=\phi(x)\setminus\{p\}$. Regarding $\phi(y)\subset\phi(x)\setminus\{p\}$: Since $y<x$, $\phi(y)\subset\phi(x)$. By construction $p\not\in\phi(y)$. Thus $\phi(y)\subset\phi(x)\setminus\{p\}$.
Regarding $\phi(x)\setminus\{p\}\subset\phi(y)$: We prove this by the contrapositive: For any $q\not\in\phi(y)$, $y\not\in q$. Define $q^{\prime}=\mathcal{L}\setminus q$, so $q^{\prime}$ is a lattice ideal and $y\in q^{\prime}$. Since $q$ is join-irreducible in $\mathcal{F}$, $q^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$. Then computing in $\mathcal{F}^{\prime}$, $q^{\prime}\geq{{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}$. Since $q^{\prime}$ is meet-irreducible and $\mathcal{F}^{\prime}$ is distributive, $q^{\prime}\geq{{x}\mathclose{\downarrow}}$ or $q^{\prime}\geq p^{\prime}$. That is equivalent to $x\in q^{\prime}$ or $q^{\prime}\geq p^{\prime}$, which is equivalent to $x\not\in q$ or $q\geq p$, which is further equivalent to $q\not\in\phi(x)$ or $q\geq p$. Since $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\not\in\phi(x)\setminus\{p\}$.
By lem. 34, $y<x$, and since $\phi(y)$ has exactly one less element than $\phi(x)$, lem. 38 shows that $y\lessdot x$. ∎
Definition 41.
Given $P\in\mathcal{I}$, we define
$D_{P}=\{Q\in\mathcal{I}\,|\,Q\mathop{\triangle}P\textup{ is a finite set}\}$
[…]
Theorem 42.
Given $P\in\mathcal{I}$, $D_{P}$ is a convex sublattice of $\mathcal{I}$.
Theorem 43.
Given $x\in\mathcal{L}$, $\phi$ is a lattice isomorphism between $\mathcal{L}$ and $D_{\phi(x)}$.
Proof.
By th. 26, $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$, so what remains to be shown is that the image of $\phi$ is $D_{\phi(x)}$.
If $y\in\mathcal{L}$, then since $\mathcal{L}$ is locally-finite, there exists a finite sequence of elements of $\mathcal{L}$, $x=z_{0},z_{1},z_{2},\ldots,z_{n}=y$ such that for every $0\leq i<n$, either $z_{i}$ covers $z_{i+1}$ or vice-versa. By lem. 36, this means that $|\phi(z_{i+1})\mathop{\triangle}\phi(z_{i})|=1$, so $|\phi(x)\mathop{\triangle}\phi(y)|\leq n$, and so $\phi(y)\in D_{\phi(x)}$.
Conversely, if $Y\in D_{\phi(x)}$, then there is a finite sequence of subsets of $\mathcal{I}$, $\phi(x)=Z_{0},Z_{1},Z_{2},\ldots,Z_{n}=Y$ such that for every $0\leq i<n$, $|Z_{i}\mathop{\triangle}Z_{i+1}|=1$. By lem. 39 and 40, this means that every $Z_{i}=\phi(z_{i})$ for some $z_{i}\in\mathcal{L}$. ∎
Thus, if we look at $\mathcal{I}$ as a graph on its elements with two elements being connected iff one covers the other, $\mathcal{L}$ is isomorphic to one connected component of $\mathcal{I}$.
In the example $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 6, $\mathcal{L}$ is isomorphic to the central connected component and there are eight other connected components. In this case, if we choose any of the eight other connected components as $\mathcal{L}^{\prime}$ and constructed the corresponding $\mathcal{I}^{\prime}$, it would be different from this $\mathcal{I}$.
However, other examples behave differently. If we choose $\mathcal{L}=\mathbb{B}_{\textup{fin}}$, the lattice of finite subsets of $\mathbb{N}$, then the corresponding $\mathcal{P}$ is isomorphic to $\mathbb{N}$, with each prime filter being all the finite sets containing some particular integer. The corresponding $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\mathbb{N}}$, the lattice of all subsets of $\mathbb{N}$. Then $\mathcal{L}$ is isomorphic to the bottom connected component of $\mathcal{I}$, containing its $\hat{0}$.
The top connected component of $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\textup{cofin}}$, the lattice of all cofinite subsets of $\mathbb{N}$, which is the lattice dual of $\mathbb{B}_{\textup{fin}}$. All other connected components are isomorphic to $\mathbb{B}_{\textup{fin}}\times\mathbb{B}_{\textup{cofin}}$. And since all components are countable but $\mathbb{B}_{\mathbb{N}}$ is uncountable, there are an uncountable number of connected components.
If we chose $\mathcal{L}$ to be any one of these components, then $\mathcal{P}$ would still be isomorphic to $\mathbb{N}$ and $\mathcal{I}$ would be the same. In this situation, for every component and every prime filter, there is some covering in the component whose separator is the prime filter.
*** Need to characterize when all prime filters are principal.
Definition 44.
[Birk1967a]*§I.7 Given two coverings $x^{-}\lessdot x^{+}$ and $y^{-}\lessdot y^{+}$ in $\mathcal{L}$, we define that $y^{-}\lessdot y^{+}$ is a downward transpose of $x^{-}\lessdot x^{+}$, in symbols $x^{+}/x^{-}\searrow y^{+}/y^{-}$, if $x^{+}=x^{-}\vee y^{+}$ and $y^{-}=x^{-}\wedge y^{+}$.
Downward transposes are transitive:
Lemma 45.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $y^{+}/y^{-}\searrow z^{+}/z^{-}$, then $x^{+}/x^{-}\searrow z^{+}/z^{-}$.
Downward transposes are directed:
Lemma 46.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $x^{+}/x^{-}\searrow z^{+}/z^{-}$, then $w^{-}\lessdot w^{+}$, $y^{+}/y^{-}\searrow w^{+}/w^{-}$, and $z^{+}/z^{-}\searrow w^{+}/w^{-}$, where $w^{-}=y^{-}\wedge z^{-}$ and $w^{+}=y^{+}\wedge z^{+}$.
Proof.
The diamond isomorphism theorem shows that $w^{-}\lessdot w^{+}$. *** ∎
Downward transposes preserve separators:
Lemma 47.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ then $x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$.
Proof.
Lemma 48.
If $x^{-}\lessdot x^{+}$ and $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}$, then
  1. (1)
    if $p$ is principal and $p={{x^{+}}\mathclose{\uparrow}}$, then there is no $x^{+}/x^{-}\searrow y^{+}/y^{-}$,
  2. (2)
    otherwise, there exists a $x^{+}/x^{-}\searrow y^{+}/y^{-}$.
Corollary 49.
If $x^{-}\lessdot x^{+}$ and $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}$, then
  1. (1)
    if $p={{y}\mathclose{\uparrow}}$, then any chain of downward transposes of $x^{-}\lessdot x^{+}$ can be extended by appending $y^{-}\lessdot y$.
  2. (2)
    if $p$ is not principal, then all chains of downward transposes of $x^{-}\lessdot x^{+}$ can be extended indefinitely.