### 15 Mar 2026 v2

# An extension of Birkhoff’s representation theorem to locally-finite distributive lattices

Abstract.
Birkhoff’s representation theorem for finite distributive lattices states that any finite distributive lattice is isomorphic to the lattice of order ideals (lower sets) of the partial order of the join-irreducible elements of the lattice. We present a simplified version of Stone’s extension of this theorem to general distributive lattices. We then apply this formulation to locally finite distributive lattices to produce a novel representation theorem: The lattice is isomorphic to the order ideals of the poset of prime filters of the lattice whose symmetric difference from a particular ideal is finite.
The 2nd version is corrected and completed.
1. Introduction
Birkhoff’s representation theorem for finite distributive lattices has long been known.
Theorem 1 (Birkhoff [Birk1937a][Birk1967a]*§III.3 p. 58).
Any finite distributive lattice $\mathcal{L}$ is isomorphic to the lattice of order ideals of the partial order of the join-irreducible elements of $\mathcal{L}$.
The essence of the proof is every element $x\in\mathcal{L}$ is mapped into the order ideal consisting of the join-irreducible elements of $\mathcal{L}$ that are $\leq x$. The the meet and join of $\mathcal{L}$ correspond to intersection and union of order ideals.
A lattice is finitary if every principal ideal (the set of elements $\leq$ a given element) is finite. Stanley shows that the representation theorem can be extended to finitary distributive lattices.
Theorem 2.
[Stan2012a]*Prop. 3.4.3 Any finitary distributive lattice $\mathcal{L}$ is isomorphic to the lattice of finite order ideals of the partial order of the join-irreducible elements of $\mathcal{L}$.
In combinatorics many interesting lattices aren’t finitary but they are locally-finite, so it is desirable to extend the representation theorem to locally-finite lattices in some way. Two obstacles immediately present themselves. One is that some locally-finite lattices have no join-irreducible elements at all, making it unclear what a representation could be constructed from. A premier example of this is $\mathbb{Z}\times\mathbb{Z}$, as shown in fig. 2.
Figure 1. The lattice $\mathbb{Z}\times\mathbb{Z}$
Figure 2. The lattice $\mathbb{Z}\times\mathbb{Z}$ with some filters outlined
The second obstacle is that in many useful cases, the lattice is isomorphic to the lattice of some, but not all, of the order ideals of join-irreducible elements. Often these ideals consist of all the ideals with the exception of the empty ideal and the ideal of all join-irreducible elements. A typical example is shown in fig. 3.
Figure 3. The lattice $t$-SkewStrip (right) and its poset of join-irreducible elements $t$-Plait (left). (Taken from [Fom1994a]*Fig. 7.)
Thus our major goal is a general theory of representation of locally-finite distributive lattices. A secondary goal is understanding lattices that are isomorphic to the lattice of some, but not all, order ideals of their join-irreducible elements.
2. Representation of non-finitary lattices
First we will lay out the analysis of general distributive lattices. Our analysis follows the discussion in [Nat2017a] and our references to it refer to chapter 8 if not otherwise specified. [^1] Prof. Nation informs me that the discussion in [Nat2017a] is derived from Bob Dilworth’s lattice theory course at Caltech in 1970–71. This discussion is an alternative to Stone’s presentation of a representation [Stone1938a] that avoids using the machinery of topology, which makes its application to discrete lattices clearer. Following the general analysis, we will apply it to locally-finite lattices.
For the remainder of this article, let $\mathcal{L}$ be a distributive lattice. As an example, consider $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 2
Definition 3.
Let $P$ be a poset. We define that $F\subset P$ is an order filter (of $P$) iff $x\in F$ and $x\leq z$ imply $z\in F$. We define that $I\subset P$ is an order ideal (of $P$) iff $x\in F$ and $z\leq x$ imply $z\in I$.
Definition 4.
[Nat2017a]*sec. 7 Let $\mathcal{L}$ be a lattice. We define that $f\subset\mathcal{L}$ is a lattice filter (of $\mathcal{L}$) iff (1) $f$ is non-empty, (2) $x,y\in f$ implies $x\wedge y\in f$, and (3) $x\in f$ and $x\leq z$ imply $z\in f$. We define that $i\subset\mathcal{L}$ is a lattice ideal (of $\mathcal{L}$) iff (1) $i$ is non-empty, (2) $x,y\in i$ implies $x\vee y\in i$, and (3) $x\in i$ and $z\leq x$ imply $z\in i$.
Definition 5.
[Nat2017a] We define that $f$ is a proper lattice filter (of a lattice $\mathcal{L}$) iff $f$ is a lattice filter and $f$ is $\neq\mathcal{L}$. We define a proper lattice ideal dually.
There is of course a natural duality on lattices, so for many facts we prove, we also state the dual fact without giving the dual proof.
Definition 6.
We define $\mathcal{F}$ to be the poset of lattice filters of $\mathcal{L}$, ordered by inverse inclusion. Dually, we define $\mathcal{F}^{\prime}$ to be the poset of lattice ideals of $\mathcal{L}$, ordered by inclusion.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, some filters (all proper) are shown in fig. 2 and the corresponding lattice $\mathcal{F}$ is shown in fig. 5.
Figure 4. The lattice of filters $\mathcal{F}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its prime elements outlined, all of which are secondary primes
Figure 5. The poset of prime filters $\mathcal{P}$ for $\mathcal{L}=\mathbb{Z}$. $\mathcal{P}=A\sqcup B$
Definition 7.
Given $f,g\in\mathcal{F}$, we define $f\mathop{\wedge\mkern-8.0mu\wedge}g=\{x\wedge y\,|\,x\in f\textup{ and }y\in g\}$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, we define $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}=\{x\vee y\,|\,x\in f^{\prime}\textup{ and }y\in g^{\prime}\}$. [^2] We pronounce $f\mathop{\wedge\mkern-8.0mu\wedge}g$ as “$f$ union-meet gg” due to lem. 9. Dually, we pronounce $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}$ as $f^{\prime}$ union-join $g^{\prime}$”.
Lemma 8.
Given $\in\mathcal{F}, fgf\mathop{\wedge\mkern-8.0mu\wedge}g$ contains $f$ and $g$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}$ contains $f^{\prime}$ and $g^{\prime}$.
Proof.
Given any $x\in f$ we can choose any $y\in g$ and then $x\vee y\in g$, so $x=x\wedge(x\vee y)$, showing $x\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Similarly, $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains $g$. ∎
Lemma 9.
Given $\in\mathcal{F}, fgf\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter and is the meet-closure of $f\cup g$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, $f^{\prime}\mathop{\wedge\mkern-8.0mu\wedge}g^{\prime}$ is a lattice ideal and is the join-closure of $f^{\prime}\cup g^{\prime}$.
Proof.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is non-empty: This is trivial because $f$ and $g$ are non-empty.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is meet-closed: Given two elements of $f\mathop{\wedge\mkern-8.0mu\wedge}g$, $x\wedge y$ and $x^{\prime}\wedge y^{\prime}$ with $x,x^{\prime}\in f$ and $y,y^{\prime}\in g$, since $f$ and $g$ are meet-closed, $x\wedge x^{\prime}\in f$ and $y\wedge y^{\prime}\in g$. Then $(x\wedge y)\wedge(x^{\prime}\wedge y^{\prime})=(x\wedge x^{\prime})\wedge(y\wedge y^{\prime})\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is closed upward: Given $x\in f$, $y\in g$ and $z\in\mathcal{L}$ with $x\wedge y\leq z$ then $z\vee x\in f$ and $z\vee y\in g$. By distributivity, $(z\vee x)\wedge(z\vee y)=z\vee(x\wedge y)=z$, showing $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$: Trivial by def. 7.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains the meet-closure of $f\cup g$: This follows from lem. 8 and the fact that $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is meet-closed. ∎
Lemma 10.
[…]
Proof.
Regarding join: That the join of two lattice filters $f$ and $g$ is $f\cap g$ is straightforward: $f\cap g$ is necessarily a lattice filter. $f\cap g$ is $\geq$ (that is, is contained in) $f$ and $g$. Clearly, any lattice filter $h\geq f,g$ is contained in both $f$ and $g$, so $h$ must be contained in $f\cap g$, and so must be $\geq f\cap g$.
Regarding meet: Given two lattice filters, $f$ and $g$, let $h=f\mathop{\wedge\mkern-8.0mu\wedge}g$. By lem. 8, $f\subset h$, which is $h\leq f$. Similarly $h\leq g$. To show that $h$ is the greatest lower bound of $f$ and $g$, assume $s$ is an lattice filter and $s\leq f,g$, that is, $f\subset s$ and $g\subset s$. Consider an arbitrary element $z$ of $h$, so $z=x\wedge y$ for some $x\in f$ and $y\in g$. Then $x$ and $y$ are $\in s$, and since $s$ is closed under meets, $z=x\wedge y\in s$. Thus $h\subset s$, which is $s\leq h$. ∎
Lemma 11.
$\mathcal{F}$ is a distributive lattice. Dually, $\mathcal{F}^{\prime}$ is a distributive lattice.
Proof.
Since we know $\mathcal{F}$ is a lattice, it suffices to prove that for any $f,g,h\in\mathcal{F}$, $f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)=(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$, which is equivalent to showing that for any $x\in\mathcal{L}$, $x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$ iff $x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$. To show that, the following are equivalent:
$x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$
$x\in f\textup{ and }x\in g\mathop{\wedge\mkern-8.0mu\wedge}h$
$x\in f\textup{ and }(\exists y\in g,z\in h)\,x=y\wedge z$ (1)
To derive (2) from (1): Since $x=y\wedge z=x\vee(y\wedge z)=(x\vee y)\wedge(x\vee z)$ and $f$, $g$, and $h$ are lattice filters,
$x\in f\textup{ and }(\exists y\in g,z\in h)\,x\vee y\in f\cap g\textup{ and }x\vee z\in f\cap h\textup{ and }x=(x\vee y)\wedge(x\vee z)$
Then choose $y^{\prime}=x\vee y$ and $z^{\prime}=x\vee z$. To derive (1) from (2): Choose $y=y^{\prime}$ and $z=z^{\prime}$, and note that $f$ is closed under meets. Thus, (1) is equivalent to:
$(\exists y^{\prime}\in f\cap g,z^{\prime}\in f\cap h)\,x=y^{\prime}\wedge z^{\prime}$ (2)
$x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$
Definition 12.
[Nat2017a] We define that a lattice filter $f$ is prime iff $f$ is proper and if $x\vee y\in f$ implies $x\in f$ or $y\in f$. We define a prime lattice ideal dually.
Lemma 13.
[Nat2017a] A lattice filter $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice ideal iff $\mathcal{L}\setminus f$ is a prime lattice ideal. Dually, a lattice ideal $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice filter iff $\mathcal{L}\setminus f$ is a prime lattice filter.
Definition 14.
For $x\in\mathcal{L}$, we define ${{x}\mathclose{\uparrow}}=\{y\in\mathcal{L}\,|\,x\leq y\}$, the principal filter generated by $x$. For $x\in\mathcal{L}$, we define ${{x}\mathclose{\downarrow}}=\{y\in\mathcal{L}\,|\,y\leq x\}$, the principal ideal generated by $x$.
Lemma 15.
[Nat2017a] For $\in\mathcal{L}, x{{x}\mathclose{\uparrow}}$ is a prime lattice filter iff $x$ is join-irreducible in $\mathcal{L}$. Dually, for $\in\mathcal{L}, x{{x}\mathclose{\downarrow}}$ is a prime lattice ideal iff $x$ is meet-irreducible in $\mathcal{L}$.
Note that if $\mathcal{L}$ were not distributive, the conditions in this theorem would have to be modified to “$x$ is join-prime” and “$x$ is meet-prime”.
Definition 16.
We define $\mathcal{P}$ to be the poset of prime elements of $\mathcal{F}$, with the order (inverse inclusion) inherited from $\mathcal{F}$. Dually, we define $\mathcal{P}^{\prime}$ to be the poset of prime elements of $\mathcal{F}^{\prime}$, with the order (inclusion) inherited from $\mathcal{F}^{\prime}$.
Lemma 17.
Lem. 13 shows that the $\omega:\mathcal{P}\rightarrow\mathcal{P}^{\prime}:p\mapsto\mathcal{L}\setminus p$ is a poset isomorphism between $\mathcal{P}$ and $\mathcal{P}^{\prime}$.
Lemma 18.
Given $x\in\mathcal{L}$ and $f\in\mathcal{F}$, the following are equivalent:
  1. (1)
    $f\leq{{x}\mathclose{\uparrow}}$,
  2. (2)
    ${{x}\mathclose{\uparrow}}\subset f$, and
  3. (3)
    $x\in f$.
Dually, given $x\in\mathcal{L}$ and $f^{\prime}\in\mathcal{F}^{\prime}$, the following are equivalent:
  1. (1)
    $f^{\prime}\leq{{x}\mathclose{\downarrow}}$,
  2. (2)
    $f^{\prime}\subset{{x}\mathclose{\downarrow}}$, and
  3. (3)
    $x\in f^{\prime}$.
Lemma 19.
[…]
In light of lem. 15 we can consider $\mathcal{F}$ to be an extension of $\mathcal{L}$ which “adds the missing join-irreducible elements of $\mathcal{L}$.”
Lemma 20.
[Nat2017a]*Th. 8.8 Sublem. A If $x\in\mathcal{L}$ and ${{x}\mathclose{\uparrow}}\not\geq f$ in $\mathcal{F}$, then there exists a $p\in\mathcal{P}$ for which $p\leq f$ but $p\not\leq{{x}\mathclose{\uparrow}}$.
Proof.
Define $\{h\in\mathcal{F}\,|\,h\leq f\textup{ and }h\not\leq{{x}\mathclose{\uparrow}}\}. HH$ is non-empty because $f\in H$. $H$ is a poset ordered by inverse inclusion, inherited from $\mathcal{F}$. Using the Hausdorff maximal principle, select $P$ which is a maximal totally ordered subset of $H$. Since $H$ is non-empty, $P$ is non-empty. Define $p=\bigcup P$. Because $P$ is a non-empty nested set of lattice filters, $\bigcup P$ is also a lattice filter, so $p\in\mathcal{F}$.
Because $x$ is in no member of $P$, $x\not\in\bigcup P=p$ and so $p\not\leq{{x}\mathclose{\uparrow}}$.
Since $P$ contains at least one element $q$, $q\leq f$, $f\subset q$, and $f\subset\bigcup P=p$, so $p\leq f$.
We prove $p$ is join-irreducible in $\mathcal{F}$: Assume there are $g,h\in\mathcal{F}$ for which $p=g\vee h$ and $g,h<p$ Since $p\not\leq{{x}\mathclose{\uparrow}}$, either $g\not\leq{{x}\mathclose{\uparrow}}$ or $h\not\leq{{x}\mathclose{\uparrow}}$. Without loss of generality, assume $g\not\leq{{x}\mathclose{\uparrow}}$, so $g\in H$.
But since $g<p=\bigcup P$, $\bigcup P\subset g$ and $\bigcup P\neq g$. Those imply that $g$ strictly contains every element of $P$ and $g$ is strictly less than (under inverse inclusion) every element of $P$, and so $g\not\in P$. Thus $P\sqcup\{g\}$ is a totally ordered subset of $\mathcal{F}$ that is strictly larger than $P$, which contradicts that $P$ is maximal. Thus $p$ is join-irreducible. ∎
The Hausdorff maximal principal is equivalent to the axiom of choice.
Corollary 21.
[Nat2017a]*Th. 8.4 Given $x\not\leq y$ in $\mathcal{L}$, there exists a $p\in\mathcal{P}$ for which $x\in p$ and $y\not\in p$.
Definition 22.
Let $p$ be an element of $\mathcal{P}$ (which thus is prime in $\mathcal{F}$). If $p={{x}\mathclose{\uparrow}}$ for some $x\in\mathcal{L}$ (which would have to be join-irreducible) we define $p$ to be a principal prime. If there is no such $x$, then we define $p$ to be a secondary prime. Both terms are relative to $\mathcal{L}$, to $\mathcal{F}$, or to $\mathcal{P}$ (by abuse of language).
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, fig. 5 shows that all primes are secondary primes. Conversely, there can be elements of $\mathcal{F}$ that are not of the form ${x}\mathclose{\uparrow}$} and are not secondary primes: $\mathbb{Z}\times\mathbb{Z}$ is in $\mathcal{F}$, is not an ${x}\mathclose{\uparrow}$}, but also is not prime.
Lemma 23.
$f\in\mathcal{F}$ is join-irreducible in $\mathcal{F}$ iff $f$ is a prime lattice filter. Dually, $f^{\prime}\in\mathcal{F}^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$ iff $f$ is a prime lattice ideal.
Proof.
We will prove the contrapositive: $f\in\mathcal{F}$ is join-reducible in $\mathcal{F}$ iff $f$ is not prime.
Regarding $\Rightarrow$: Assume $f$ is join-reducible. Then there exist filters $g$ and $h$ for which $g,h\neq f$ and $f=g\vee h=g\cap h$. Choose $x\in g\setminus f$ and $y\in h\setminus f$; thus $x,y\not\in f$. Since $g$ and $h$ are filters, $x\vee y\in g,h$, and so $x\vee y\in f$. Thus $f$ is not prime.
Regarding $\Leftarrow$: Assume $f$ is not prime. Then there exists $x$ and $y$ with $x,y\not\in f$ but $x\vee y\in f$. Define $g={{x}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$ and $h={{y}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$, both of which are $\neq f$. We will show that $f=g\vee h=g\cap f$. By lem. 8, $f\subset g\cap h$. Conversely, suppose $z\in g\cap h$. Then there exists $u,u^{\prime},v,v^{\prime}\in f$ for which $z=u\wedge u^{\prime}=v\wedge v^{\prime}$, $u\geq x$, $v\geq y$, and $u^{\prime},v^{\prime}\in f$. Then $z=(u\wedge u^{\prime})\vee(v\wedge v^{\prime})=(u\vee v)\wedge(u\vee v^{\prime})\wedge(u^{\prime}\vee v)\wedge(u^{\prime}\vee v^{\prime})$. But all four terms of that meet are $\in f$, so $z\in f$, showing $g\cap h\subset f$. Thus, $f=g\cap h=g\vee h$, showing $f$ is join-reducible. ∎
Definition 24.
We define $\mathcal{I}$ to be the lattice of order ideals of $\mathcal{P}$ under inclusion, which has $\cap$ as meet and $\cup$ as join. Dually, $\mathcal{I}$ is lattice-isomorphic to the lattice of order ideals of $\mathcal{P}^{\prime}$.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, the corresponding lattice $\mathcal{I}$ is shown in fig. 6.
Figure 6. The lattice of ideals $\mathcal{I}$ of $\mathcal{P}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its connected components outlined.
Definition 25.
We define $\phi(x)=\{p\in\mathcal{P}\,|\,x\in p\}$.
Lem. 18 can be extended:
Lemma 26.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$, the following are equivalent:
  1. (1)
    $p\leq{{x}\mathclose{\uparrow}}$,
  2. (2)
    ${{x}\mathclose{\uparrow}}\subset p$,
  3. (3)
    $x\in p$, and
  4. (4)
    $p\in\phi(x)$.
Theorem 27.
[Nat2017a]*Th. 8.5 $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$.
Corollary 28.
Every distributive lattice can be represented as a “ring of sets” under intersection and union.
3. Application to finitary lattices
Applying the general theory to finitary lattices is straightforward.
Theorem 29.
If $\mathcal{L}$ is finitary then every element $p\in\mathcal{F}$ is ${x}\mathclose{\uparrow}$} for some $x\in\mathcal{L}$.
Proof.
Let $p\in\mathcal{F}$. Since $p$ is non-empty and $\mathcal{L}$ is finitary, there must be a minimal element $x\in p$. Since $p$ is closed under meets, every element of $y\in p$ must be $\geq x$, as otherwise $y\wedge x$ would be be <x<x and $\in p$. Thus $x$ is minimum in $p$. Since $p$ is a lattice filter, $p={{x}\mathclose{\uparrow}}$. ∎
Corollary 30.
[Stan2012a]*Prop. 3.4.3 If $\mathcal{L}$ is finitary then: (1) $\mathcal{P}$ contains no secondary primes, so $\mathcal{P}$ is (isomorphic to) the poset of join-irreducible elements of $\mathcal{L}$. (2) Every value of $\phi$ is a finite set in $\mathcal{I}$. (3) For every finite $I\in\mathcal{I}$, $\phi(\bigvee I)=I$. (4) $\phi$ is a lattice isomorphism between $\mathcal{L}$ and the set of finite elements of $\mathcal{I}$.
4. Application to locally-finite lattices
In this section, we apply the general theory to lattices that are locally-finite but not necessarily finitary. Henceforth we assume that $\mathcal{L}$ is locally-finite.
Lemma 31.
If $x\in\mathcal{L}$ is join-irreducible, then $x$ covers exactly one element of $\mathcal{L}$. Dually, if $x\in\mathcal{L}$ is meet-irreducible, then $x$ is covered by exactly one element of $\mathcal{L}$. (The converses of these statements are trivial.)
Proof.
Since $x$ is join-irreducible, it cannot be the $\hat{0}$ of $\mathcal{L}$ if there is one, so there is an element $y<x$. Since $\mathcal{L}$ is locally-finite, there must be an element $z$ such that $y\leq z\lessdot x$. If $x$ covered more than one element of $\mathcal{L}$, then $x$ would be the join of any two of those elements and would not be join-irreducible. ∎
Lemma 32.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$ where $x\not\in p$, There exists a $y\in\mathcal{L}$ for which $y>x$, $y\in p$, and $y$ is the minimum element of ${{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$ and thus ${{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$.
Proof.
Define $f={{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$. Since $f$ is a lattice filter, it is not empty, and we can choose $y\in f$, so $y\geq x$. Since $x\not\in f$, $y>x$. And since $\mathcal{L}$ is locally-finite and all choices for $y$ are bounded below by $x$, we can choose $y$ to be minimal in $f$. Any minimal $y$ must be the minimum of $f$, because if $y_{1}\neq y_{2}$ are both minimal in $f$, since $f$ is a lattice filter, $y_{1}\wedge y_{2}$ is also in $f$, contradicting that $y_{1}$ and $y_{2}$ are minimal in $f$. Thus, $f={{y}\mathclose{\uparrow}}$. ∎
Definition 33.
Given $x\in\mathcal{L}$ and $p\in\mathcal{P}$ where $x\not\in p$, we define $x\mathord{\Uparrow}p$ to be the minimum element of ${{x}\mathclose{\uparrow}}\vee p={{x}\mathclose{\uparrow}}\cap p$.
We define the dual operation on lattice ideals:
Lemma 34.
Given $x\in\mathcal{L}$ and $p^{\prime}\in\mathcal{P}^{\prime}$ where $x\not\in p^{\prime}$, There exists a $y\in\mathcal{L}$ for which $y<x$, $y\in p^{\prime}$, and $y$ is the maximum element of ${{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\uparrow}}\cap p^{\prime}$ and thus ${{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\downarrow}}\cap p^{\prime}$.
Proof.
Define $f^{\prime}={{x}\mathclose{\downarrow}}\wedge p^{\prime}={{x}\mathclose{\downarrow}}\cap p^{\prime}$. Since $f^{\prime}$ is a lattice ideal, it is not empty, and we can choose $y\in f^{\prime}$, so $y\leq x$. Since $x\not\in f^{\prime}$, $y<x$. And since $\mathcal{L}$ is locally-finite and all choices for $y$ are bounded above by $x$, we can choose $y$ to be maximal in $f^{\prime}$. Any maximal $y$ must be the minimum of $f^{\prime}$, because if $y_{1}\neq y_{2}$ are both maximal in $f$, since $f^{\prime}$ is a lattice ideal, $y_{1}\vee y_{2}$ is also in $f^{\prime}$, contradicting that $y_{1}$ and $y_{2}$ are maximal in $f^{\prime}$. Thus, $f^{\prime}={{y}\mathclose{\downarrow}}$. ∎
Definition 35.
Given $x\in\mathcal{L}$ and $p^{\prime}\in\mathcal{P}^{\prime}$ where $x\not\in p^{\prime}$, we define $x\mathord{\Downarrow}p^{\prime}$ to be the maximum element of ${{x}\mathclose{\downarrow}}\wedge p={{x}\mathclose{\downarrow}}\cap p^{\prime}$.
Lemma 36.
If $x\lessdot y$ in $\mathcal{L}$, then $\phi(x)\subset\phi(y)$ and $\phi(y)\setminus\phi(x)$ contains exactly one element.
Proof.
By def. 25, it is clear that $\phi(x)\subset\phi(y)$.
By cor. 21, there exists a $p\in\mathcal{P}$ for which $y\in p$ and $x\not\in p$. Thus $p\in\phi(y)\setminus\phi(x)$.
Assume that $p\neq q$ are both $\in\phi(y)\setminus\phi(x)$. Consider $\mathord{\Uparrow}p: x<x$$x\mathord{\Uparrow}p\leq y$. Since $x\lessdot y$, $x\mathord{\Uparrow}p=y$ and $p={{y}\mathclose{\uparrow}}$. Similarly $q={{y}\mathclose{\uparrow}}$, and so $p=q$.
Thus $\phi(y)\setminus\phi(x)$ has exactly one element. ∎
Definition 37.
Given a covering $x\lessdot y$, we define the separator of $x$ and $y$, $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y$ to be the unique element of $\phi(y)\setminus\phi(x)$. (Thus $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y\in\mathcal{P}$.)
Lemma 38.
$\mathcal{L}$ is graded. If $x\leq y$ in $\mathcal{L}$, then $\phi(x)\subset\phi(y)$ and $|\phi(y)\setminus\phi(x)|=\rho(x,y)$, where $\rho(x,y)$ is the rank difference of $y$ above $x$.
Lemma 39.
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\not\in\phi(x)$ and $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$. Then $x\lessdot x\mathord{\Uparrow}p$ and $\phi(x\mathord{\Uparrow}p)=\phi(x)\sqcup\{p\}$.
Proof.
Define $y=x\mathord{\Uparrow}p$.
We will show that $\phi(y)=\phi(x)\sqcup\{p\}$. Regarding $\phi(x)\sqcup\{p\}\subset\phi(y)$: Since $x<y$, $\phi(x)\subset\phi(y)$. Since $y\in p$, $p\in\phi(y)$. Thus $\phi(x)\sqcup\{p\}\subset\phi(y)$.
Regarding $\phi(y)\subset\phi(x)\sqcup\{p\}$: For any $q\in\phi(y)$, $q\leq{{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p$. Since $\mathcal{P}$ is distributive and $q$ is join-irreducible in $\mathcal{P}$ by lem. 23, $q\leq{{x}\mathclose{\uparrow}}$ or $q\leq p$. That is equivalent to $q\in\phi(x)$ or $q\leq p$. Since $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\in\phi(x)\sqcup\{p\}$. Thus $\phi(y)\subset\phi(x)\sqcup\{p\}$.
Together, these show $\phi(y)=\phi(x)\sqcup\{p\}$. By lem. 32, $x<y$, and since $\phi(y)$ has exactly one more element than $\phi(x)$, lem. 38 shows that $x\lessdot y$. ∎
Lemma 40.
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\in\phi(x)$ and $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$. Then defining $p^{\prime}=\mathcal{L}\setminus p$, $x\gtrdot x\mathord{\Downarrow}p^{\prime}$ and $\phi(x\mathord{\Downarrow}p^{\prime})=\phi(x)\setminus\{p\}$.
Proof.
(This proof is dual to the proof of lem. 39 but that is concealed by our notation, since $\phi$ is not self-dual.)
Define $p^{\prime}=\mathcal{L}\setminus p$ and $y=x\mathord{\Downarrow}p^{\prime}$.
We will show that $\phi(y)=\phi(x)\setminus\{p\}$. Regarding $\phi(y)\subset\phi(x)\setminus\{p\}$: Since $y<x$, $\phi(y)\subset\phi(x)$. By construction $p\not\in\phi(y)$. Thus $\phi(y)\subset\phi(x)\setminus\{p\}$.
Regarding $\phi(x)\setminus\{p\}\subset\phi(y)$: We prove this by the contrapositive: For any $q\not\in\phi(y)$, $y\not\in q$. Define $q^{\prime}=\mathcal{L}\setminus q$, so $q^{\prime}$ is a lattice ideal and $y\in q^{\prime}$. Since $q$ is join-irreducible in $\mathcal{F}$ by lem. 23, $q^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$. Then computing in $\mathcal{F}^{\prime}$, $q^{\prime}\geq{{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}$. Since $q^{\prime}$ is meet-irreducible and $\mathcal{F}^{\prime}$ is distributive, $q^{\prime}\geq{{x}\mathclose{\downarrow}}$ or $q^{\prime}\geq p^{\prime}$. That is equivalent to $x\in q^{\prime}$ or $q^{\prime}\geq p^{\prime}$, which is equivalent to $x\not\in q$ or $q\geq p$, which is further equivalent to $q\not\in\phi(x)$ or $q\geq p$. Since $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\not\in\phi(x)\setminus\{p\}$.
By lem. 34, $y<x$, and since $\phi(y)$ has exactly one less element than $\phi(x)$, lem. 38 shows that $y\lessdot x$. ∎
Definition 41.
Given $P\in\mathcal{I}$, we define
$D_{P}=\{Q\in\mathcal{I}\,|\,Q\mathop{\triangle}P\textup{ is a finite set}\}$
[…]
Lemma 42.
Given $P\in\mathcal{I}$, $D_{P}$ is a convex sublattice of $\mathcal{I}$.
Theorem 43.
Given $x\in\mathcal{L}$, $\phi$ is a lattice isomorphism between $\mathcal{L}$ and $D_{\phi(x)}$.
Proof.
By th. 27, $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$, so what remains to be shown is that the image of $\phi$ is $D_{\phi(x)}$.
If $y\in\mathcal{L}$, then since $\mathcal{L}$ is locally-finite, there exists a finite sequence of elements of $\mathcal{L}$, $x=z_{0},z_{1},z_{2},\ldots,z_{n}=y$ such that for every $0\leq i<n$, either $z_{i}$ covers $z_{i+1}$ or vice-versa. By lem. 36, this means that $|\phi(z_{i+1})\mathop{\triangle}\phi(z_{i})|=1$, so $|\phi(x)\mathop{\triangle}\phi(y)|\leq n$, and so $\phi(y)\in D_{\phi(x)}$.
Conversely, if $Y\in D_{\phi(x)}$, then there is a finite sequence of subsets of $\mathcal{I}$, $\phi(x)=Z_{0},Z_{1},Z_{2},\ldots,Z_{n}=Y$ such that for every $0\leq i<n$, $|Z_{i}\mathop{\triangle}Z_{i+1}|=1$. By lem. 39 and 40, this means that every $Z_{i}=\phi(z_{i})$ for some $z_{i}\in\mathcal{L}$. ∎
Thus, if we look at $\mathcal{I}$ as a graph whose vertices are the elements of $\mathcal{I}$ with two elements being connected iff one covers the other, $\mathcal{L}$ is isomorphic to one connected component of $\mathcal{I}$.
In the example $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 6, $\mathcal{L}$ is isomorphic to the central connected component, and there are eight other connected components. In this case, if we chose any of the eight other connected components as $\mathcal{L}_{1}$, the corresponding $\mathcal{P}_{1}$ would be a proper subset of $\mathcal{P}$, and $\mathcal{I}_{1}$ would be a proper subset of this $\mathcal{I}$.
However, other examples behave differently. If we choose $\mathcal{L}=\mathbb{B}_{\textup{fin}}$, the lattice of finite subsets of $\mathbb{N}$, then the corresponding $\mathcal{P}$ is an antichain with $\aleph_{0}$ elements, with each prime filter being all the finite sets containing some particular integer and each prime filter is a principal prime. The corresponding $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\mathbb{N}}$, the lattice of all subsets of $\mathbb{N}$. Then $\mathcal{L}$ is isomorphic to the bottom connected component of $\mathcal{I}$, containing its $\hat{0}$$\emptyset$.
The top connected component of $\mathcal{I}$ contains its $\hat{1}=\mathcal{P}$ and is isomorphic to $\mathbb{B}_{\textup{cofin}}$, the lattice of all cofinite subsets of $\mathbb{N}$, which is the lattice dual of $\mathbb{B}_{\textup{fin}}$. All other connected components are isomorphic to $\mathbb{B}_{\textup{fin}}\times\mathbb{B}_{\textup{cofin}}$. And since all components are countable but $\mathbb{B}_{\mathbb{N}}$ is uncountable, there are an uncountable number of connected components of $\mathcal{I}$.
If we chose $\mathcal{L}_{1}$ to be any one of these components other than $\mathbb{B}_{\textup{fin}}$, then $\mathcal{P}_{1}$ would be an antichain with $\aleph_{0}$ elements, but its elements would all be secondary primes. The corresponding $\mathcal{I}_{1}$ would be isomorphic to $\mathcal{I}$.
5. Criterion for secondary primes
Given a locally-finite lattice, if all of its primes are principal, they correspond to the join-irreducible elements of $\mathcal{L}$, and so the structure of $\mathcal{P}$ is relatively easy to determine. But if a prime is secondary, it has no single corresponding element of $\mathcal{L}$ and so determining the structure of $\mathcal{P}$ is more difficult. We will now show how to distinguish these two cases.
Definition 44.
[Birk1967a]*§I.7 Given two coverings $x^{-}\lessdot x^{+}$ and $y^{-}\lessdot y^{+}$ in $\mathcal{L}$, we define that $y^{-}\lessdot y^{+}$ is a downward transpose of $x^{-}\lessdot x^{+}$, in symbols $x^{+}/x^{-}\searrow y^{+}/y^{-}$, if $x^{+}=x^{-}\vee y^{+}$ and $y^{-}=x^{-}\wedge y^{+}$.
Downward transposes are transitive:
Lemma 45.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $y^{+}/y^{-}\searrow z^{+}/z^{-}$, then $x^{+}/x^{-}\searrow z^{+}/z^{-}$.
Downward transposes are directed:
Lemma 46.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $x^{+}/x^{-}\searrow z^{+}/z^{-}$, then defining $w^{-}=y^{-}\wedge z^{-}$ and $^{+}=y^{+}\wedge z^{+}, ww$$^{-}\lessdot w^{+}$, $y^{+}/y^{-}\searrow w^{+}/w^{-}$, and $z^{+}/z^{-}\searrow w^{+}/w^{-}$.
Proof.
By hypothesis, $y^{-}=y^{+}\wedge x^{-}$, $z^{-}=z^{+}\wedge x^{-}$, $x^{+}=x^{-}\vee z^{+}$, and $y^{+}\leq x^{+}$. Then $y^{-}\wedge w^{+}=y^{-}\wedge(y^{+}\wedge z^{+})=(y^{+}\wedge x^{-})\wedge(y^{+}\wedge z^{+})=(y^{+}\wedge x^{-})\wedge(z^{+}\wedge x^{-})=y^{-}\wedge z^{-}=w^{-}$ and $y^{-}\vee w^{+}=(y^{+}\wedge x^{-})\vee(y^{+}\wedge z^{+})=y^{+}\wedge(x^{-}\vee z^{+})=y^{+}\wedge x^{+}=y^{+}$. Together, these show $y^{+}/y^{-}\searrow w^{+}/w^{-}$. Similarly, we show $z^{+}/z^{-}\searrow w^{+}/w^{-}$. Applying the diamond transposition theorem to either downward transpose shows that $w^{-}\lessdot w^{+}$. ∎
Downward transposes preserve separators (def. 37):
Lemma 47.
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ then $x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$.
Proof.
We know that $\phi(x^{+})\setminus\phi(x^{-})$ and $\phi(y^{+})\setminus\phi(y^{-})$ each contain exactly one element. Define $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}=\phi(x^{+})\setminus\phi(x^{-})$, so $p\not\in\phi(x^{-})$, $x^{-}\not\in p$, and $x^{-}\vee y^{+}=x^{+}\in p$. Since $p$ is a prime filter, either $x^{-}\in p$ or $y^{+}\in p$. But by hypothesis, $x^{-}\not\in p$, so $y^{+}\in p$. Since $x^{-}\not\in p$ by hypothesis and $y^{-}\leq x^{-}$, $y^{-}\not\in p$. Thus $y\in\phi(y^{+})\setminus\phi(y^{-})$, so $p=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$. ∎
Theorem 48.
If $x^{-}\lessdot x^{+}$ and $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}$, then
  1. (1)
    if $p$ is principal and $p={{x^{+}}\mathclose{\uparrow}}$, then there is no $x^{+}/x^{-}\searrow y^{+}/y^{-}$ with $x^{+}>y^{+}$.
  2. (2)
    otherwise, there exists a $y^{-}\lessdot y^{+}$ with $x^{+}>y^{+}$ and $x^{+}/x^{-}\searrow y^{+}/y^{-}$.
Proof.
Regarding (1): If there was such a $y^{-}\lessdot y^{+}$, then by lem. 47, $p=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$ and $y^{+}\in p$. But that would mean that $y^{+}\geq x^{+}$, but by hypothesis $x^{+}>y^{+}$.
Regarding (2): If $p$ is not principal, or $p\neq{{x^{+}}\mathclose{\uparrow}}$, then there exists a $y^{+}\lessdot x^{+}$, and necessarily $y^{+}\neq x^{-}$. Define $y^{-}=y^{+}\wedge x^{-}$. Then by the diamond isomorphism theorem, $y^{-}\lessdot y^{+}$ and $x^{+}/x^{-}\searrow y^{+}/y^{-}$. ∎
Definition 49.
We define a chain of downward transposes to be a sequence $x^{-}_{1}\lessdot x^{+}_{1},x^{-}_{2}\lessdot x^{+}_{2},x^{-}_{3}\lessdot x^{+}_{3},\ldots,x^{-}_{n}\lessdot x^{+}_{n}$ with $x^{+}_{1}>x^{+}_{2}>x^{+}_{3}>\cdots>x^{+}_{n}$ (or equivalently $x^{-}_{1}>x^{-}_{2}>x^{-}_{3}>\cdots>x^{-}_{n}$) and $x_{1}^{+}/x_{1}^{-}\searrow x_{2}^{+}/x_{2}^{-}$, $x_{2}^{+}/x_{2}^{-}\searrow x_{3}^{+}/x_{3}^{-}$, $\ldots$, $x_{n-1}^{+}/x_{n-1}^{-}\searrow x_{n}^{+}/x_{n}^{-}$.
Corollary 50.
If $x^{-}\lessdot x^{+}$ and $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}$, then
  1. (1)
    if $p$ is principal, then any chain of downward transposes of starting with $x^{-}\lessdot x^{+}$ can only be extended to some finite maximum length.
  2. (2)
    if $p$ is secondary, then all chains of downward transposes starting with $x^{-}\lessdot x^{+}$ can be extended indefinitely.
Cor. 50 can be used to discern which prime filters are primary and which are secondary. For example, in fig. 7(a) all chains of downward transposes can be extended indefinitely, so all primes are secondary.
But in fig. 7(b) all chains of downward transposes must end, showing all primes are principal. The upper elements of the coverings on which chains are forced to end are the join-irreducible elements of $\mathcal{L}$, and collectively they are are isomorphic to $\mathcal{P}$. $\mathcal{I}$ has three connected components, one is $\emptyset$, one is $\mathcal{P}$, and one is the remaining ideals of $\mathcal{P}$, all of which are infinite and have complements that are infinite. The third connected component is isomorphic to this lattice.
(a) The lattice $\mathbb{Z}\times\mathbb{Z}$ with a chain of downward transposes
(b) The lattice 44-SkewStrip with several chains of downward transposes
Figure 7.
6. Unique connected component of ideals
The examples we have been using, $\mathbb{Z}\times\mathbb{Z}$ and $\mathbb{B}_{\textup{fin}}$, share the property that their lattice of ideals $\mathcal{I}$ has multiple connected components, one of which is isomorphic the the lattice that generates it. However, our long-term goal is characterizing lattices which are useful in constructing RSK algorithms. Examples of these lattices are $t$-SkewStrip in fig. 3 and the lattices listed in [Wor2026b]*sec. 2.2. All of those lattices have the property that that their lattice of ideals $\mathcal{I}$ has at most three components:
  • one component that is isomorphic to the generating lattice $\mathcal{L}$;
  • if $\mathcal{L}$ has no $\hat{1}$, one component containing only $\mathcal{P}$; and
  • if $\mathcal{L}$ has no $\hat{0}$, one component containing only $\emptyset$.
Looking at these examples suggests:
Conjecture 51.
A locally-finite distributive lattice has a lattice of ideals of prime filters that contains at most three connected components (one isomorphic to the lattice, perhaps one containing the empty ideal, and perhaps one containing the ideal of all primes) iff the lattice has finite width (contains no infinite antichains).